OpenStudy (anonymous):

Wirte an equation for the line that passes through (-1,-1) and (1,-5).

OpenStudy (zpupster):

slope intercept formula y=mx+b to find the slope m=y2-y1/x2-x1

OpenStudy (anonymous):

Since m=-2, what does b equal?

OpenStudy (zpupster):

if you let x be 0 and solve you b-(or your y-intercept) is that

OpenStudy (anonymous):

Since m=-2, you have a point to test, and you know the equation \[y=mx+b\] plug in a point you have (let's use (1,-5) this time), and the m you found and solve for b:\[\rightarrow b=y-mx=(-5)-(-2)(1)=-3\]

OpenStudy (anonymous):

So, the final answer is y=-2x-3?

OpenStudy (anonymous):

Yes, that's correct. You can verify this by putting the equation in a graphic calculator and going to the table. You'll find that both the points (-1,-1) and (1,-5) are part of the same line.

OpenStudy (anonymous):

Thanks